# Dynamics of Meta-learning Representation in the Teacher-student Scenario

### v1

@@ Appendix C Details on the Variability of Teacher Representation Mapping
In this section, we provide more details of the case where the teacher networks of different tasks have different representation mapping as introduced in Sec. IV.5 of the main text. In this case, the $t$-th task produces samples as
$\sigma^{t}=\frac{1}{\sqrt{M}}\sum_{m=1}^{M}u^{t}_{m}\;g\big{(}(\gamma\boldsymbol{B}_{m}+\sqrt{1-\gamma^{2}}\Delta\boldsymbol{B}_{m}^{t})^{\top}\cdot\boldsymbol{\xi}\big{)}.$ (53)
@@
$\sigma^{t}\approx\frac{1}{\sqrt{M}}\sum_{m=1}^{M}u_{m}^{t}\;\big{[}g(\gamma\boldsymbol{B}_{m}^{\top}\cdot\boldsymbol{\xi})+\sqrt{1-\gamma^{2}}g^{\prime}(\gamma\boldsymbol{B}_{m}^{\top}\cdot\boldsymbol{\xi})(\Delta\boldsymbol{B}_{m}^{t\top}\cdot\boldsymbol{\xi})\big{]}$ (54)
@@
Appendix D L2-norm Regularization
In this section, we enrich the modeling of meta-learning by considering two additional elements : (i) the outputs are noisy, (ii) L2-norm regularization for the representation mapping $\mathbf{J}$ is used. For the latter case, it can help to stabilize the solution of the meta-learner’s weights $\mathbf{J}$.
@@ Appendix D L2-norm Regularization
So, we can update the parameter $\boldsymbol{J}_{k}$ as
@@
$-2\lambda\eta_{J}Q_{kl}.$ (67)
Where $I_{2}^{\prime}(k,l)$ can claculate as
@@
Appendix E Linear Activation Function
E.1 Dynamical Equations
Throughout the paper, we have analyzed meta-learning on networks with nonlinear activation functions. Here, we turn our attention to two-layer networks with a linear activation function $g(x)=x$ as considered in [17, 18]. In this case, it is possible for each hidden unit of the meta-learner to learn a superposition of all hidden units of the meta-teachers, so we need to examine the overlap order parameters more carefully to determine weather the teacher’s representation mappings $\mathbf{B}$ have been properly learned.
@@ E.1 Dynamical Equations
Following the same procedures outlined in the main text, we obtain the following dynamical equations for the order parameters and the expression for the meta-generalization errors as following
@@ E.2 Case Study
In the second choice of initial conditions, the results in Fig. 11(d) show that $\rho_{11}=\rho_{21}=\rho_{31}=0$, while the other $\rho_{kn}$ values are non-zero and differ from $1$. This implies that $\boldsymbol{J}_{1}$, $\boldsymbol{J}_{2}$, and $\boldsymbol{J}_{3}$ have overlaps with $\boldsymbol{B}_{2}$ and $\boldsymbol{B}_{3}$ but remain orthogonal to $\boldsymbol{B}_{1}$. As shown in Fig. 11(b), $Q_{22}=Q_{33}\neq Q_{11}$, $Q_{12}=Q_{13}=0$, and $Q_{23}\neq 0$, indicating that $\boldsymbol{J}_{1}$ is orthogonal to $\boldsymbol{J}_{2}$ and $\boldsymbol{J}_{3}$. Therefore, in this case, the meta-learner has learned a two-dimensional sub-space of the meta-teacher’s representation mapping.
@@
Finally, in the third choice of initial conditions, the results in Fig. 12(d) show that all $\rho_{kn}$ values are non-zero and distinct from one. Fig. 12(b) shows that $Q_{11}=Q_{22}=Q_{33}\neq 0$ and $Q_{kl}=0$ for $k\neq l$, indicating that $\boldsymbol{J}_{1}$, $\boldsymbol{J}_{2}$, and $\boldsymbol{J}_{3}$ are mutually orthogonal. We can deduce that the linear space spanned by $\{\boldsymbol{J}_{1},\boldsymbol{J}_{2},\boldsymbol{J}_{3}\}$ coincides with the meta-teacher network’s representation space spanned by $\{\boldsymbol{B}_{1},\boldsymbol{B}_{2},\boldsymbol{B}_{3}\}$. Therefore, in this case, the student network has fully learned the teacher network’s representation, and as a result, the meta-generalization error is able to decrease to zero as shown in Fig. 12(c). These results are consistent with the findings in Refs. [17] and [18]. The results also suggest that a higher degree of heterogeneity in the initial conditions, especially for $\{R_{kn}\}$, is needed for effective meta-learning in linear networks.
@@
Figure 10: Evolution of the meta-generalization error and the quantities $Q$, $R$, and $\rho$ in the case with linear activation function, with the system parameters specified in Sec. E.2. The first set of initial condition for $\{R_{kn}\}$ is used, i.e., $R_{kn}=1\times 10^{-12}$ at $\alpha=0$. (a) The evolution of the meta-generalization error. (b) $Q_{kl}$ v.s. $\alpha$. (c) $R_{kn}$ v.s. $\alpha$. (d) $\rho_{kn}$ v.s. $\alpha$.
@@
Figure 11: Evolution of the meta-generalization error and the quantities $Q$, $R$, and $\rho$ in the case with linear activation function, with the system parameters specified in Sec. E.2. The second set of initial condition for $\{R_{kn}\}$ is used, i.e., $R_{11}=1.1\times 10^{-12}$ with the remaining elements set as $R_{kn}=1\times 10^{-12}$ at $\alpha=0$. (a) The evolution of the meta-generalization error. (b) $Q_{kl}$ v.s. $\alpha$. (c) $R_{kn}$ v.s. $\alpha$. (d) $\rho_{kn}$ v.s. $\alpha$.
@@
Figure 12: Evolution of the meta-generalization error and the quantities $Q$, $R$, and $\rho$ in the case with linear activation function, with the system parameters specified in Sec. E.2. The third set of initial condition for $\{R_{kn}\}$ is used, i.e., $R_{11}=1.1\times 10^{-12}$, $R_{31}=1.2\times 10^{-12}$, with the remaining elements set as $R_{kl}=1\times 10^{-12}$ at $\alpha=0$. (a) The evolution of the meta-generalization error. (b) $Q_{kl}$ v.s. $\alpha$. (c) $R_{kn}$ v.s. $\alpha$. (d) $\rho_{kn}$ v.s. $\alpha$.

### v2

@@ Appendix D Details on the Variability of Teacher Representation Mapping
In this appendix, we provide more details of the case where the teacher networks of different tasks have different representation mapping as introduced in Sec. IV.5. In this case, the $t$th task produces samples as
$\sigma^{t}=\frac{1}{\sqrt{M}}\sum_{m=1}^{M}u^{t}_{m}\;g\big{[}(\gamma\boldsymbol{B}_{m}+\sqrt{1-\gamma^{2}}\Delta\boldsymbol{B}_{m}^{t})^{\top}\cdot\boldsymbol{\xi}\big{]}.$ (59)
@@
$\sigma^{t}\approx\frac{1}{\sqrt{M}}\sum_{m=1}^{M}u_{m}^{t}\;\big{[}g(\gamma\boldsymbol{B}_{m}^{\top}\cdot\boldsymbol{\xi})+\sqrt{1-\gamma^{2}}g^{\prime}(\gamma\boldsymbol{B}_{m}^{\top}\cdot\boldsymbol{\xi})(\Delta\boldsymbol{B}_{m}^{t\top}\cdot\boldsymbol{\xi})\big{]}.$ (60)
@@
Appendix E L2-norm Regularization
In this appendix, we enrich the modeling of meta-learning by considering two additional elements: (i) the outputs are noisy and (ii) L2-norm regularization for the representation mapping $\mathbf{J}$ is used. For the latter case, it can help to stabilize the solution of the meta-learner’s weights $\mathbf{J}$.
@@ Appendix E L2-norm Regularization
So we can update the parameter $\boldsymbol{J}_{k}$ as
@@
$-2\lambda\eta_{J}Q_{kl},$ (73)
where $I_{2}^{\prime}(k,l)$ can be calculated as
@@
Appendix F Linear Activation Function
F.1 Dynamical equations
Throughout the paper, we have analyzed meta-learning on networks with nonlinear activation functions. Here we turn our attention to two-layer networks with a linear activation function $g(x)=x$ as considered in [17, 18]. In this case, it is possible for each hidden unit of the meta-learner to learn a superposition of all hidden units of the meta-teachers, so we need to examine the overlap order parameters more carefully to determine weather the teacher’s representation mappings $\mathbf{B}$ have been properly learned.
@@ F.1 Dynamical equations
Following the same procedures outlined in the main text, we obtain the following dynamical equations for the order parameters and the expression for the meta-generalization errors:
@@ F.2 Case study
In the second choice of initial conditions, the results in Fig. 11(d) show that $\rho_{11}=\rho_{21}=\rho_{31}=0$, while the other $\rho_{kn}$ values are nonzero and differ from $1$. This implies that $\boldsymbol{J}_{1}$, $\boldsymbol{J}_{2}$, and $\boldsymbol{J}_{3}$ have overlaps with $\boldsymbol{B}_{2}$ and $\boldsymbol{B}_{3}$ but remain orthogonal to $\boldsymbol{B}_{1}$. As shown in Fig. 11(b), $Q_{22}=Q_{33}\neq Q_{11}$, $Q_{12}=Q_{13}=0$, and $Q_{23}\neq 0$, indicating that $\boldsymbol{J}_{1}$ is orthogonal to $\boldsymbol{J}_{2}$ and $\boldsymbol{J}_{3}$. Therefore, in this case, the meta-learner has learned a two-dimensional subspace of the meta-teacher’s representation mapping.
@@
Finally, in the third choice of initial conditions, the results in Figure 12(d) show that all $\rho_{kn}$ values are non-zero and distinct from one. Figure 12(b) shows that $Q_{11}=Q_{22}=Q_{33}\neq 0$ and $Q_{kl}=0$ for $k\neq l$, indicating that $\boldsymbol{J}_{1}$, $\boldsymbol{J}_{2}$, and $\boldsymbol{J}_{3}$ are mutually orthogonal. We can deduce that the linear space spanned by $\{\boldsymbol{J}_{1},\boldsymbol{J}_{2},\boldsymbol{J}_{3}\}$ coincides with the meta-teacher network’s representation space spanned by $\{\boldsymbol{B}_{1},\boldsymbol{B}_{2},\boldsymbol{B}_{3}\}$. Therefore, in this case, the student network has fully learned the teacher network’s representation, and as a result, the meta-generalization error is able to decrease to zero as shown in Fig. 12(c). These results are consistent with the findings in Refs. [17] and [18]. The results also suggest that a higher degree of heterogeneity in the initial conditions, especially for $\{R_{kn}\}$, is needed for effective meta-learning in linear networks.
@@
Figure 10: Evolution of the meta-generalization error and the quantities $Q$, $R$, and $\rho$ in the case with linear activation function, with the system parameters specified in Sec. F.2. The first set of initial condition for $\{R_{kn}\}$ is used, i.e., $R_{kn}=1\times 10^{-12}$ at $\alpha=0$. (a) The evolution of the meta-generalization error. (b) $Q_{kl}$ vs $\alpha$. (c) $R_{kn}$ vs $\alpha$. (d) $\rho_{kn}$ vs $\alpha$.
@@
Figure 11: Evolution of the meta-generalization error and the quantities $Q$, $R$, and $\rho$ in the case with linear activation function, with the system parameters specified in Sec. F.2. The second set of initial condition for $\{R_{kn}\}$ is used, i.e., $R_{11}=1.1\times 10^{-12}$ with the remaining elements set as $R_{kn}=1\times 10^{-12}$ at $\alpha=0$. (a) The evolution of the meta-generalization error. (b) $Q_{kl}$ vs $\alpha$. (c) $R_{kn}$ vs $\alpha$. (d) $\rho_{kn}$ vs $\alpha$.
@@
Figure 12: Evolution of the meta-generalization error and the quantities $Q$, $R$, and $\rho$ in the case with linear activation function, with the system parameters specified in Sec. F.2. The third set of initial condition for $\{R_{kn}\}$ is used, i.e., $R_{11}=1.1\times 10^{-12}$, $R_{31}=1.2\times 10^{-12}$, with the remaining elements set as $R_{kl}=1\times 10^{-12}$ at $\alpha=0$. (a) The evolution of the meta-generalization error. (b) $Q_{kl}$ vs $\alpha$. (c) $R_{kn}$ vs $\alpha$. (d) $\rho_{kn}$ vs $\alpha$.